OpenStudy (anonymous):

Write a polynomial equation of the smallest degree with roots -i, i and -2.

OpenStudy (anonymous):

(x^2+1)(x+2)

OpenStudy (anonymous):

start with \((x+i)(x-i)(x+2)\) and multiply out

OpenStudy (anonymous):

umm would i start like this \[p(x) = (x + i) (x - i) (x + 2)\]

OpenStudy (anonymous):

you get as a first step what @ChmE wrote,then multiply those together

OpenStudy (anonymous):

FOIL them right?

OpenStudy (anonymous):

That is a good starting point. I just skipped a step. Yes foil them

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i got \[x^3 + 2x^2 = x^2(x + 2)\]